An automobile radiator holds four gallons and contains 25% antifreeze. How much of the solution must be drained from the radiator and antifreeze added to make the solution test 35% antifreeze?
Answer by stanbon(75887)

An automobile radiator holds four gallons and contains 25% antifreeze. How much of the solution must be drained from the radiator and antifreeze added to make the solution test 35% antifreeze?
--------------------------
Equation:
anti - anti + anti = anti
--------
0.25*4 - 0.25*x + 1.00x = 0.35*4
---------
25*4 - 25x + 100x = 35*4
-----
75x = 40
x = 8/15 gallon (amt. of pure antifreeze to add)
=============
Cheers,
Stan H.
=============